OpenStudy (anonymous):

what is 230000000000 in scientific notaion

OpenStudy (jdoe0001):

\(\Large 23\underline{0000000000}\implies 23\times 10^{\textit{how many zeros?}}\)

OpenStudy (anonymous):

So in scientific notation, you only keep the first number in the ones place, then all the other numbers will contribute to the exponent on the 10. lets start off with our original: 230000000000 is the same thing as 230000000000 * 10^0 since 10^0 = 1 now every time we divide our original by 10, to get rid of a place, we have to add 1 to the exponent for example, 230000000000 = 23000000000.0 * 10^1 = 2300000000.00 * 10^2 the beauty of this is that for very large numbers or very small decimals with lots of zeros, you can save yourself the time of writing all the zeros (normally we wouldnt include the numbers after the decimal if they were zeros). in this case we have 10 zeros and a 3 that we want to put behind the decimal Thats 11 number places. That means we can rewrite our number with a 10^11 230000000000 = 2.3 * 10^11

OpenStudy (anonymous):

@jdoe0001 I believe you should have a decimal there :)

OpenStudy (jdoe0001):

we.... you can always add it I guess